OpenStudy (anonymous):

6(x-11)/5 = 5(x+11)/6 So get 36(x-11) = 25 * (x +11) How is it getting the 36 and the 25? I can solve this problem after figuring those two out. But, I am not getting how to get those two numbers in the first place.

OpenStudy (anonymous):

\[\frac{ 6(x-11) }{ 5 }=\frac{ 5(x+11) }{ 6 }\]You're going to cross multiply, but first you can distribute you're 6(x-11) and 5(x+11):\[\frac{ 6x-66 }{ 5 }=\frac{ 5x+55 }{ 6 }\]Now, cross multiply: 5(5x+55)=6(6x-66), then distribute: 25x+275=36x-396